# Infinite-Horizon Optimal Wireless Control Over Shared State-Dependent Fading Channels for IIoT Systems

### v1

@@ Abstract
Heterogeneous systems consisting of a multiloop wireless control system (WCS) and a mobile agent system (MAS) are ubiquitous in Industrial Internet of Things systems. Within these systems, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over and can significantly compromise its performance. This paper focuses on the infinite-horizon optimal control of MAS to ensure the performance of WCS while minimizing an average cost for the heterogeneous system subject to state and input constraints. Firstly, the state-dependent fading channel is modeled, which characterizes the interference among transmission links, and shows that the probability of a successful transmission for WCS depends on the state of MAS. A necessary and sufficient condition in terms of constrained set stabilization is then established to ensure the Lyapunov-like performance of WCS with expected decay rate. Secondly, using the semi-tensor product of matrices and constrained reachable sets, a criterion is presented to check the constrained set stabilization of MAS and to ensure the performance of WCS. In addition, a constrained optimal state transition graph is constructed to address state and input constraints, by resorting to which the feasibility analysis of the optimal control problem is presented. Finally, an algorithm is proposed for the construction of optimal input sequences by minimum-mean cycles for weighted graph. An illustrative example is provided to demonstrate effectiveness of the proposed method.
@@
Constrained set stabilization; heterogeneous Industrial Internet of Things system; infinite-horizon optimal control; shared wireless medium.
@@ 1.1 Background and motivation
Industrial Internet of Things (IIoT) enables flexible, efficient and sustainable production in many fields including smart manufacturing through numerous plants that can be reconfigured based on process requirements [Baumann2021, fadlullah2011]. These plants consist of multiple physical systems with sensing, computation, communication and actuation capabilities, which flexibly interact with one another and human operators [Ahlen2019, vitturi2019]. Due to such complex interactions among the entities, wired communication will reach its limits. Instead, wireless sensors are deployed to monitor industrial plants, which enable much higher flexibility while reducing cable wrapping and limitations on plant operations, forming wireless control systems (WCSs). This abundance of wireless devices spawns the need of efficiently sharing the available wireless medium [gatsis2018].
@@
An integral aspect of IIoT in smart manufacturing is the use of mobile agents [Pulikottil2021], which can coordinate with WCSs contributing to overall manufacturing processes in various ways, forming a heterogeneous IIoT system, e.g., manufacturing systems with heavy machines and cranes [Agrawal2014], assembly processes with autonomous assembly arms and forklifts [hu2022, hu2019]. In these scenarios, the state of agents are subject to certain constraints to ensure production safety such as collision and dangerous position avoidance. Moreover, positions of mobile agents may induce shadow fading on the wireless channel that the WCS is controlled over, leading to variations in the measured channel gains [Agrawal2014, Quevedo2013]. In order to ensure the performance of WCSs, it is necessary to explicitly examine the influence of mobile agents’ positions on channel states of the WCS, and design constrained agents motion strategy such that the desired performance requirements is guaranteed for the WCS. Such coupling brings heterogeneity to IIoT systems, namely, the WCS is coupled with the mobile agent system (MAS), which is a distinctive feature of IIoT systems compared with traditional WCSs.
Some results considered the dependence of channel state on physical states, which mainly focused on channel characterization, and communication and controller design of WCSs [Agrawal2014, Ahlen2019, Kashiwagi2010]. Few works addressed the heterogeneity feature induced by this dependence. A state-dependent channel model was proposed in [hu2019] to characterize the WCS and MAS coupling, and approximate optimal control and transmission power policies were obtained, under which the WCS achieves both safety and efficiency. This work was generalized in [hu2022] by considering a state-dependent Markov channel model to address time varying data rates, where optimal strategies were obtained by solving a constrained polynomial optimization problem. In these results, the state and input constraints for the MAS were ignored. In addition, only a simplified model consisting of a WCS with a single control loop and a single agent was considered. How to solve the coordination problem of WCSs and MASs in IIoTs with multiple plants being controlled over a shared wireless channel is largely unexplored. The main challenge is to model the fading channel which captures both the interference among transmission links, and the coupling between WCS and MAS, based on which to address the heterogeneity of the whole IIoT system.
@@ 1.2 Related work
As a part of IIoT, WCSs with fading channels have been extensively studied in the past few decades. It is noted that the channel gains used to characterize shadow fading are traditionally modeled as independent identical distributed random processes [Gatsis2014] or Markov chains [zhangq1999], where the network state is assumed to be independent from physical states. With such independency, several results were established for WCSs over fading channels including co-design problem for communication and control systems [Gatsis2014], state estimation [Quevedo2012] and so on. References [gatsis2015] and [gatsis2018] considered shared fading channels, where centralized and distributed channel-aware access mechanisms were respectively presented to meet the Lyapunov-like performance requirements while minimizing the total expected power consumption. It is obvious that these methods on performance analysis of WCS with fading channels fail in the IIoT scenario where the WCS and MAS coupling exists, not to mention the case of constrained MAS.
@@
Few works addressing dependence of network state on physical states have been conducted in IIoT. A state-dependent bursty fading channel model was proposed in [hu2013], based on which distributed switching control [hu2015] and power control schemes [hu2020] were respectively proposed to ensure the performance of vehicular WCSs. However, only homogeneous systems without constraints were considered in these works. In order to model random time variations of wireless environment, [Quevedo2013] introduced a network state process for state estimation, but the network topology was assumed to be fixed. The optimal network topology configuration was then determined for each network state to minimize an expected error covariance measure [Leong2016]. However, the external environment in [Leong2016, Quevedo2013] was modeled as a (semi)-Markov chain which cannot be controlled. Removing the uncontrollability assumptions and modeling the external environment as a Markov decision process, [hu2019] studied the infinite-horizon optimal design problem of communication and control, such that the single-loop WCS with state-dependent channel achieves both safety and efficiency. By formulating this problem as a constrained two-player cooperative game, approximate optimal control and transmission power policies were obtained by solving relaxed convex generalized geometric programs. Similar problem was studied in [hu2022], which considered a generalized state-dependent Markov channel model to address time varying data rates. By solving a constrained polynomial optimization problem, optimal design strategies were obtained. In the above works, safety constraints of the mobile agents were ignored. In addition, for the practical smart manufacturing scenarios where several plants are controlled over a shared wireless medium, and coordinate with multiple mobile agents to complete manufacturing tasks, the proposed state-dependent channel models cannot capture the interference among transmission links, and the coupling between WCS and MAS.
As for the other part of IIoT, recently, the finite field network architecture, which is a special kind of finite-valued systems, has been adopted to model the analysis and operational control for MAS [Pasqualetti2014]. The (in)finite horizon optimal control problems for finite-valued systems have received extensive attention [Fornasini2014, wuy2019, Zhao2011]. It is noted that most existing results are established for the optimal control of finite-valued systems with only system dynamics constraints. Taking the state constraints into account, [gao2021-1] proposed an efficient graph-theoretical approach for switched finite-valued systems. The optimal problem with a set stabilization constraint was considered in [gao2021], where only the finite-horizon optimization until reaching the target set was addressed. Noting that all the above results are established for homogeneous finite-valued systems, i.e., the dynamics of all nodes in the system is the same, they are not applicable for the optimization of heterogeneous IIoT systems subject to performance requirements constraints.
@@ 1.3 Contributions
In this paper, we focus on the optimal control of MAS with state and input constraints to ensure the Lyapunov-like performance of WCS, while minimizing an infinite-horizon average cost for the IIoT system in the presence of WCS and MAS coupling. A preliminary version of the paper appears in [wangs2022], where only controller design problem is addressed for the MAS to ensure the performance of WCS. The main contributions of the paper are summarized as follows.
@@
First, to take into account the level of shadowing effects on channel gains of the WCS, the finite field network architecture is used to model the operational control for the MAS. The state-dependent fading channel is then modeled to characterize the dependence of wireless transmission of the WCS on mobile agents’ positions, leading to the state-dependent Lyapunov-like performance requirement of the WCS. Based on this model, the infinite-horizon optimal control problem is established for the heterogeneous IIoT system to quantify the tradeoff between WCS’s power consumption and MAS’s control cost, while ensuring the Lyapunov-like performance requirement of the WCS and the safety constraints of the MAS.
Second, optimal input sequences are designed for the MAS to solve the infinite-horizon optimal control problem. A criterion in terms of constrained set stabilization for the MAS is established to ensure the Lyapunov-like performance of WCS, which equivalently converts the optimal control problem into the infinite-horizon optimal constrained set stabilization problem of the MAS. Since each agent updates its states from a finite field, it is difficult to apply classical nonlinear system theory to solve the optimal problem. We adopt the semi-tensor product approach to represent the MAS as a discrete-time bilinear system, based on which a necessary and sufficient condition is presented for the constrained set stabilization of MAS. In order to address state and input constraints of the MAS, a constrained optimal state transition graph is constructed, by analyzing which the feasibility of the considered optimal control problem is proved. Then, an algorithm is proposed to design optimal input sequences by resorting to minimum-mean cycles in weighted graph.
@@
This paper is organized as follows. In Section 2, we characterize the heterogeneous IIoT system and formulate the problem studied in this paper. Section 3 considers Lyapunov-like performance analysis of the WCS. Feasibility analysis and optimal controller design are proceeded in Section 4 and Section 5, respectively. Section 6 demonstrates main results by an illustrative example, which is followed by the conclusion in Section 7. The key notations used in this paper are summarized in Table 1. Throughout this paper, semi-tensor product is the basic matrix product defined as $M\ltimes P=(M\otimes I_{l/n})(P\otimes I_{l/p})$, where $M\in{\mathbb{R}}^{m\times n}$, $P\in{\mathbb{R}}^{p\times q}$, $l$ is the least common multiple of $n$ and $p$ [chengqi2011, cheng2011]. Since the semi-tensor product is a generalization of the conventional matrix product, the symbol $\ltimes$” is omitted in most places of this paper when no confusion arises.
@@
$\mathcal{D}_{n}^{m}$ Cartesian product $\underbrace{\mathcal{D}_{n}\times\cdots\times\mathcal{D}_{n}}_{m}$
@@
$\delta_{s}[i_{1}\;i_{2}\;\cdots\;i_{t}]$ Logical matrix $F$ with $[F]_{:,j}=\delta_{s}^{i_{j}}$
${\mathcal{L}}^{n\times m}$ Set of $n\times m$ logical matrices
@@
$\otimes$ Kronecker product
@@
$\mathcal{M}\setminus\mathcal{N}$ Set $\{x\in\mathcal{M}:x\notin\mathcal{N}\}$
@@ 2 System model and problem formulation
Figure 1: Framework of a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a factory floor, discretized into $\kappa$ subregions. ‘S’ and ‘A’ respectively represent sensing and actuation capabilities. The local channel state of each control loop $i$ ($\gamma_{i}\in\mathcal{D}_{r}$) is dependent on positions of agents $\alpha_{j}\in\mathcal{D}_{\kappa}$, $j=1,\cdots,n$.
We consider a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a typical smart manufacturing scenario (Fig. 1). In the wireless control architecture, $q$ independent plants are controlled over a shared wireless medium, where the measurements of plant $i$, $i\in\{1,\cdots,q\}$ are wirelessly transmitted to the access point by sensor $i$ to compute the control inputs. At the same time, $n$ mobile agents perform manufacturing tasks in the same workspace. In this process, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over, resulting in the coupling between the WCS and MAS. According to the level of shadowing effects on channel gains of the WCS, partition all possible positions of the mobile agents in the factory floor into $\kappa$ overlapping sets. Each set corresponds to a certain level of shadowing effects. Correspondingly, the whole workspace is discretized into $\kappa$ overlapping planar subregions, called cells $\{0,1,\cdots,\kappa-1\}$.
@@ 2.1 Plants
Denote the state of plant $i$ at time $k$ by $x_{i}(k)\in\mathbb{R}^{n_{i}\times 1}$, $i=1,\cdots,q$. The discrete-time model of plant $i$ is
$x_{i}(k+1)=A_{i}x_{i}(k)+B_{i}v_{i}(k)+\xi_{i}(k),$ (1)
where $A_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $B_{i}\in\mathbb{R}^{n_{i}\times m_{i}}$ are system matrix and input matrix, respectively, $v_{i}\in\mathbb{R}^{m_{i}\times 1}$ is the driving input, and $\{\xi_{i}(k):k\in\mathbb{N}\}$ represents an independent identically distributed noise process with mean zero and real positive semi-definite covariance $\Xi_{i}$.
@@ 2.2 Sensors and access point
$q$ sensors are placed to monitor the plants. At every slot $k$, each sensor $i$ takes a measurement and chooses whether transmitting over the shared channel or not adapting to its local wireless channel state. We adopt the discrete channel model [Sadeghi2008], i.e., the local channel state of system $i$ at slot $k$ is $\gamma_{i}(k)\in\mathcal{D}_{r}$. Assume that the wireless communication policy is pre-designed as
$\varphi_{i}(k)=h_{i}(\gamma_{i}(k)),$ (2)
where $h_{i}:\mathcal{D}_{r}\rightarrow\mathcal{D}_{2}$, $\varphi_{i}(k)=1$ (or 00) means that sensor $i$ transmits (or does not transmit) over the shared channel at slot $k$. In practice, before deciding whether to transmit over the shared channel, each sensor can acquire the channel state information by either directly using the visual sensor to observe the positions of the mobile agents, or a short pilot signal sent from the access point to the sensor at the beginning of each time slot. The pilot signal may also serve as a synchronization purpose [gatsis2018, hu2011].
When sensor $i$ decides to transmit, it wirelessly transmits measurements of plant $i$ to the access point with a fixed power $\mu_{i}>0$, where the access point plays the role of remote controller responsible for computing control decisions based on the received measurements [ma2018]. Then, control decisions are wirelessly fed back to actuator $i$. These $q$ independent control loops form a multiloop WCS.

### v2

@@ Abstract
Heterogeneous systems consisting of a multiloop wireless control system (WCS) and a mobile agent system (MAS) are ubiquitous in Industrial Internet of Things systems. Within these systems, the positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over and can significantly compromise its performance, requiring joint coordination between the WCS and MAS. Such coordination introduces different time steps and hybrid state spaces consisting of logical components and continuous components. This paper focuses on the infinite-horizon optimal control of MAS to ensure the performance of WCS while minimizing an average cost for the heterogeneous system subject to safety constraints. A state-dependent fading channel is modeled to capture interference among transmission links, as well as the effects of mobile agents’ movements on successful wireless transmission. In order to address the heterogeneous system dynamics, the optimal control problem is formulated as the optimal constrained set stabilization of the MAS by establishing a necessary and sufficient condition for the Lyapunov-like performance of WCS with the expected decay rates. Using the semi-tensor product of matrices, a constrained optimal state transition graph is constructed to encode the constrained system dynamics as well as objective function, which further reduces the problem into a minimum-mean cycle problem for the graph. By studying the properties of the graph, the feasibility is proven, and an effective algorithm is proposed for the construction of optimal input sequences. An illustrative example is provided to demonstrate effectiveness of the proposed method.
@@
Heterogeneous system; infinite-horizon optimal control; shared wireless medium; constrained set stabilization; semi-tensor product of matrices.
@@ 1.1 Background and motivation
Industrial Internet of Things (IIoT) enables flexible, efficient and sustainable production in many fields, including smart manufacturing, through numerous plants that can be reconfigured based on process requirements [Baumann2021, fadlullah2011]. These plants consist of multiple physical systems with sensing, computation, communication and actuation capabilities, which flexibly interact with one another and human operators [Ahlen2019, vitturi2019]. Due to such complex interactions among the entities, wired communication will reach its limits. Instead, wireless sensors are deployed to monitor industrial plants, which enable much higher flexibility while reducing cable wrapping and limitations on plant operations, forming multiloop wireless control systems (WCSs).
@@
An integral aspect of IIoT in smart manufacturing is the use of mobile agents, e.g., automated guided vehicles (AGVs) [Pulikottil2021], which can coordinate with WCSs contributing to overall manufacturing processes in various ways, e.g., manufacturing systems with heavy machines and cranes [Agrawal2014], assembly processes with autonomous assembly arms and forklifts [hu2022, hu2019]. In these scenarios, the state of agents are subject to certain constraints to ensure production safety such as moving within task areas and velocity limits while avoiding dangerous positions [Baumann2021]. Moreover, the positions of mobile agents may induce shadow fading on the wireless channel that the WCS is controlled over [Agrawal2014, Quevedo2013]. On-site measurements in factories reveal that the channel fading caused by machinery movements exhibits a piecewise behavior, i.e., the fading channel gain remains stationary as the machinery moves within a certain area [Ahlen2019]. And each area is associated with a certain level of shadowing effects. This corresponds to a high level dynamics of MAS exhibiting logical characteristics. Additionally, the movements of mobile agents among areas after spatial discretization is typically on the time scale of seconds, while the wireless channel transmits packets on the order of milliseconds. This means that the coordination of the WCS and mobile agent system (MAS) introduces different time steps and hybrid state spaces, consisting of the continuous component of industrial plants and the logical component of mobile agents. These distinctive features make the modeling and analysis tools on traditional WCSs fail in addressing the IIoT system considered in the paper. Due to the complex interactions between the WCS and MAS, in order to ensure the control performance of the WCS, it is necessary to explicitly examine the influence of mobile agents’ positions on the channel states of the WCS, and to design a constrained motion strategy for agents such that the desired performance requirements are guaranteed for the WCS.
@@ 1.2 Related work
As important components of IIoT systems, the MAS and WCS are typically considered separately in the literature. Recently, the finite-field network architecture, which is a special kind of finite-valued systems, has been adopted to model the analysis and operational control for MAS [Pasqualetti2014]. The (in)finite horizon optimal control problems for finite-valued systems have received extensive attention [Fornasini2014, wuy2019, Zhao2011]. It is noted that most existing results are established for the optimal control of finite-valued systems with only system dynamics constraints. Taking the state constraints into account, [gao2021-1] proposed an efficient graph-theoretical approach for switched finite-valued systems. Noting that all the above results are established for finite-valued systems, i.e., the dynamics of all nodes in the system are the same, they are not applicable for the optimization of heterogeneous IIoT systems subject to control performance requirements constraints. As for the other part of IIoT, WCSs with fading channels have been extensively studied in the past few decades. It is noted that the channel gains used to characterize shadow fading are traditionally modeled as independent identical distributed random processes [Gatsis2014] or Markov chains [zhangq1999], where the network state is assumed to be independent from physical states. With such independency, several results were established for WCSs over fading channels including co-design problem for communication and control systems [Gatsis2014], state estimation [Quevedo2012] and so on. References [gatsis2015] and [gatsis2018] considered shared fading channels, where centralized and distributed channel-aware access mechanisms were respectively presented to meet the Lyapunov-like performance requirements while minimizing the total expected power consumption. The above channel models are clearly inadequate for characterizing the complex interactions between the WCS and MAS. In addition, all these results focus on the control performance of plants with only continuous state space. It is obvious that these methods fail in analyzing IIoT systems with hybrid state spaces considered in this paper.
@@
Few works addressing dependence of network state on physical states have been conducted in IIoT. A state-dependent bursty fading channel model was proposed in [hu2014]. However, only homogeneous systems without constraints were considered in these works. In order to model random time variations of wireless environment, [Quevedo2013] introduced a network state process for state estimation, but the network topology was assumed to be fixed. The optimal network topology configuration was then determined for each network state to minimize an expected error covariance measure [Leong2016]. However, the external environment in [Leong2016, Quevedo2013] was modeled as a (semi)-Markov chain which cannot be controlled. Removing the uncontrollability assumptions and modeling the external environment as a Markov decision process, [hu2019] studied the infinite-horizon optimal design problem of communication and control, such that the single-loop WCS with state-dependent channel achieves both safety and efficiency. By formulating this problem as a constrained two-player cooperative game, approximate optimal control and transmission power policies were obtained by solving relaxed convex generalized geometric programs. Similar problem was studied in [hu2022], which considered a generalized state-dependent Markov channel model to address time varying data rates. By solving a constrained polynomial optimization problem, optimal design strategies were obtained. In the above works, a common time step is adopted for the evolution of the MAS and WCS. Noting that the wireless channel typically transmit packets on the order of milliseconds, this requires mobile agents to move on a millisecond time scale, which is often unrealistic. In addition, for the practical smart manufacturing scenarios where several plants are controlled over a shared wireless medium, the state-dependent channel models cannot capture the interference among transmission links, and the coupling between WCS and MAS.
@@ 1.3 Contributions
In this paper, we focus on the optimal control of MAS with safety constraints to ensure the Lyapunov-like performance of WCS, while minimizing an infinite-horizon average cost for the IIoT system in the presence of WCS and MAS coupling. A preliminary version of the paper appears in [wangs2022], where a common time step is considered for the WCS and the leader-follower MAS. The main contributions of the paper are summarized as follows.
@@
  • We propose a heterogeneous IIoT system framework to capture the coupling between the WCS and MAS, where the multiloop WCS is described as a switched dynamics model depending on wireless transmission, and the finite-field architecture is used to model the high level dynamics of MAS moving among areas with different levels of shadow effects. Different time steps are adopted to capture different time scales between the wireless transmission of control systems and the movements of mobile agents among discretized areas. Effect of mobile agents’ movements on successful wireless transmission is characterized by a state-dependent fading channel, which captures the spectrum resource competition and interference among different loops. Based on this model, the infinite-horizon optimal control problem is established for the heterogeneous IIoT system to quantify the tradeoff between WCS’s power consumption and MAS’s control cost, while ensuring the Lyapunov-like performance requirement of the WCS and the safety constraints of the MAS.
  • To address the hybrid nature of the heterogeneous IIoT system, we first equivalently convert the optimal control problem into the optimal constrained set stabilization problem of the MAS. In order to address the nonlinearity due to the finite-field model of the MAS, we adopt the semi-tensor product approach, and construct a constrained optimal state transition graph to handle safety constraints elegantly. By encoding the constrained system dynamics and objective function into the graph, the problem is further deduced into a minimum-mean cycle problem on the constrained optimal state transition graph. Then, through studying the properties of the graph, we propose an efficient algorithm to design optimal input sequences that ensure wireless control performance and safety constraints.
@@
This paper is organized as follows. In Section 2, we characterize the heterogeneous IIoT system and formulate the problem studied in this paper. Section 3 formulates the optimal control problem as the optimal constrained set stabilization problem for the MAS. Design of optimal controllers using the graphical method is provided in Section 4. Section 5 demonstrates main results by an illustrative example, which is followed by the conclusion in Section 6. The key notations used in this paper are summarized in Table 1. Throughout this paper, semi-tensor product is the basic matrix product defined as $M\ltimes P=(M\otimes I_{l/n})(P\otimes I_{l/p})$, where $M\in{\mathbb{R}}^{m\times n}$, $P\in{\mathbb{R}}^{p\times q}$, $l$ is the least common multiple of $n$ and $p$, and $\otimes$ is the Kronecker product of matrices [chengqi2011, cheng2011]. Since the semi-tensor product is a generalization of the conventional matrix product, the symbol $\ltimes$” is omitted in most places of this paper when no confusion arises.
@@
[…]
${}\mathbb{P}\{\cdot\}$ Probability of a random event
$\mathbb{E}\{\cdot\}$ Expectation of a random variable
@@ 2 System model and problem formulation
We consider a heterogeneous IIoT system where a multiloop WCS and an MAS coordinate with each other to jointly perform overall tasks in a typical smart manufacturing scenario (Fig. 1). In the wireless control architecture, $q$ independent plants are controlled over a shared wireless medium, where the measurements of plant $i$, $i\in\{1,\cdots,q\}$ are wirelessly transmitted to the access point by sensor $i$ to compute the control inputs. At the same time, $n$ mobile agents perform manufacturing tasks in the same workspace. In this process, positions of mobile agents may lead to shadow fading on the wireless channel that the WCS is controlled over, resulting in the coupling between the WCS and MAS.
Figure 1: Framework of a heterogeneous IIoT consisting of a multiloop WCS and an MAS in a factory floor. ‘S’ and ‘A’ respectively represent sensing and actuation capabilities. The workspace is partitioned into $\kappa$ cells $\{0,1,\cdots,\kappa-1\}$, and each cell is abstracted as a logical state indicating a level of shadowing effects. Each mobile agent $j$ moves within its task area, leading to logical state transitions among cells in the state constraint set $\mathcal{C}_{\alpha_{j}}$.
@@ 2.1 Plants
Denote the state of plant $i$ at time instant $t_{l}$ by $x_{i}(l)\in\mathbb{R}^{n_{i}}$, $i=1,\cdots,q$. The discrete-time model of plant $i$ is
$x_{i}(l+1)=A_{i}x_{i}(l)+B_{i}v_{i}(l)+\xi_{i}(l),$ (1)
where $\{t_{l}\}_{l=0}^{\infty}$ with $t_{l}<t_{l+1}$ is the sequence of sampling time instants, $A_{i}\in\mathbb{R}^{n_{i}\times n_{i}}$ and $B_{i}\in\mathbb{R}^{n_{i}\times m_{i}}$ are system matrix and input matrix, respectively, $v_{i}\in\mathbb{R}^{m_{i}}$ is the driving input, and $\{\xi_{i}(l):l=0,1,\cdots\}$ represents an independent identically distributed noise process with mean zero and real positive semi-definite covariance $\Xi_{i}$.
@@ 2.2 Sensors and access point
$q$ sensors are placed to monitor the plants. At every time instant $t_{l}$, each sensor $i$ takes a measurement and chooses whether transmitting over the shared channel or not. We adopt the discrete channel model [Sadeghi2008], i.e., the local channel state of system $i$ at time instant $t_{l}$ is $\gamma_{i}(l)\in\mathcal{D}_{r}$. Assume that the wireless communication policy111 Regardless of the effect of mobile agents’ movements on wireless transmission, following the approach in [gatsis2018], it can be proven that the optimal transmission policies ensuring the Lyapunov-like control performance of the WCS only depend on the channel states $\gamma_{i}(l)$. is pre-designed as
$\varphi_{i}(l)=h_{i}(\gamma_{i}(l)),$ (2)
where $h_{i}:\mathcal{D}_{r}\rightarrow\mathcal{D}_{2}$, $\varphi_{i}(l)=1$ (or 00) means that sensor $i$ transmits (or does not transmit) over the shared channel at time instant $t_{l}$, and each sensor transmits only when its corresponding local channel state is favorable enough and backs off otherwise. In practice, before deciding whether to transmit over the shared channel, each sensor can acquire the local channel state information by either directly using the visual sensor to observe the positions of the mobile agents, or a short pilot signal sent from the access point to the sensor at the beginning of each time slot. The pilot signal may also serve as a synchronization purpose [gatsis2018, hu2011].
When sensor $i$ decides to transmit, it wirelessly transmits measurements of plant $i$ to the access point with a fixed power $\mu_{i}>0$, where the access point plays the role of remote controller responsible for computing control decisions based on the received measurements [ma2018]. Then, control decisions are wirelessly fed back to actuator $i$. These $q$ control loops form a multiloop WCS.